OpenStudy (anonymous):

\[\lim_{n \to \infty} \frac{x^n}{n!} \text{ for } x >1\]

OpenStudy (anonymous):

a hint would be perfect :-D

OpenStudy (jamesj):

do you know the answer?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

n! i believe is domineering

OpenStudy (anonymous):

log lele

OpenStudy (anonymous):

x.x.x.x.../1.2.3.4...

OpenStudy (anonymous):

hmm differentiation wrt n?

OpenStudy (anonymous):

take log

OpenStudy (anonymous):

i am :-D him

OpenStudy (jamesj):

sorry, no l'hopital. Instead, let N = [|x|] + 1. Then for every n > N |x/n| < 1. Using this you can show the product must go to zero.

OpenStudy (amistre64):

\[\frac{x^{n+1}}{(n+1)!}*\frac{n!}{x^n}\] \[\frac{x^{n+1}}{(n+1)*n!}*\frac{n!}{x^n}\] \[\frac{x}{(n+1)}\] n+1 takes the lead no matter what x you can think of

OpenStudy (anonymous):

but why'd you take N=[|x|] +1james

OpenStudy (jamesj):

i.e., the n-th term for n > N is \[ |a_n| = \frac{|x|^N}{N!} \frac{|x|^{n-N}}{n!/(N-n)!} < \frac{|x|^N}{N!} \frac{|x|^{n-N}}{N^{n-N}} \]

OpenStudy (jamesj):

Now, show the the RHS --> 0, which it will because |x/N| < 1. That is why we chose N = [|x|]+1

OpenStudy (jamesj):

correction: that middle fraction should be \[ \frac{|x|^{n-N}}{n!/N!} \]

OpenStudy (jamesj):

Now, knowing that for |a| < 1, \[ \lim_{n \rightarrow \infty} a^n = 0 \] it follows that \[ \lim_{n \rightarrow \infty} |x/N|^{n-N} = 0 \]

OpenStudy (anonymous):

I need to work on it, thanks btw